OpenStudy (anonymous):

WRITE THE EQUIVALENT RATE ~ 15/1h =900ft/1min how did i get this show work

OpenStudy (anonymous):

Show work !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

hellooooooooooooooooooo

OpenStudy (anonymous):

huh?????????

jimthompson5910 (jim_thompson5910):

there's a typo when you typed "15/1h =900ft/1min " what's after the first 15?

OpenStudy (anonymous):

well this is the equation....

jimthompson5910 (jim_thompson5910):

is it ft? miles?

jimthompson5910 (jim_thompson5910):

or something else

OpenStudy (anonymous):

15mi/1h =900ft/1h.

jimthompson5910 (jim_thompson5910):

ok that changes everything

OpenStudy (anonymous):

i need to show how i got this

OpenStudy (anonymous):

900ft/ 1hr or 900ft/ 1 min??

jimthompson5910 (jim_thompson5910):

1 mile = 5280 ft (this is just something to be looked up or memorized) multiply both sides by 15 to get 15 miles = 79200 ft So 15 mi/1 hr = 79200 ft/1 hr ------------------------------------ Now, 1 hr = 60 min So replace 1 hr in the denominator with 60 min to get 79200 ft/1 hr = 79200 ft/60 min Now divide each component by 60 to get 79200/ 60 = 1320 and 60/60 = 1 This means that 79200 ft/60 min = 1320 ft/1 min So the answer is 1320 ft/1 min Make sure you typed everything correctly

OpenStudy (anonymous):

^^ yeah thats what i got too.

jimthompson5910 (jim_thompson5910):

so either there's a typo in the problem or a typo in the answer they gave of 900 ft/1 min